You wish to test the following claim at a significance level of α=0.05

      H0:p=0.65
      H1:p≠0.65

You obtain a sample of size n=684 in which there are 436 successful observations.
What is the test statistic for this sample? (Report answer accurate to 3 decimal places.)
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p =0.65

Ha : p 0.65

= x / n = 436/684=0.637

P0 = 0.65

1 - P0 = 1-0.65=0.35

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

=0.637-0.65 / [(0.65*0.35) / 684]

=-0.713